# Learning Task Representations from In-Context Learning

### v1

@@ C.3.2 Prompting
[…]
@@ C.3.3 Evaluation
We uniformly sample in-context examples (for few-shot) and a query input $N=256$ times, assessing performance as the accuracy if the first token generated by the model matches the first token of the query label (if the label consists of more than one token). To compute the margin of error (denoted after $\pm$ in Table 1), we first calculate the standard error of the mean (SEM) as follows:
SEM=standard $\operatorname{SEM}=\frac{\text{standard deviation}}{\sqrt{N}}.$
Next, the t-critical value is determined using the t-distribution with a 95% confidence level and degrees of freedom ($N-1$). The margin of error is then calculated as:
Margin of $\text{Margin of error}=t_{\text{critical}}\times\text{SEM}.$
@@ C.4.1 Learning from Regression Data
We recognize that generating distinct functions and prompts at each gradient step could potentially provide an infinite variety of data and functions. This raises concerns about whether LTV overfits to the function classes. Although it was argued that the likelihood of the model encountering a training-similar prompt is extremely low [8], we opted for a static dataset approach. Our experiments were conducted using this dataset, with evaluations performed on the prompts constructed from different weight vectors during inference.
@@ C.6 Detailed Ablation Studies
We estimate the probability distributions using Kernel Density Estimation (KDE), employing the standard KDE implementation from scipy [52] with default settings. However, directly estimating a probability distribution in a high-dimensional space often leads to the curse of dimensionality, where the volume of data required to effectively estimate the distribution grows exponentially with the number of dimensions. A practical solution to this challenge is to employ SVD for dimensionality reduction. This involves decomposing the data matrices $X_{i}\in\mathbb{R}^{M\times d}$ as:
@@ C.7 Computational Resources
Appendix D Complementary Results
In this section, we present the complete set of results that could not be accommodated in the main body due to space constraints. Specifically, we include comprehensive evaluations of the regression data, results under distributional shifts, and histograms for the ablation studies.
D.1 Complete Set of Evaluations on Regression Tasks
In addition to the main text, we present results where the LTV is trained with $T_{\mathbf{v}}=T_{\text{train}}$ and $T_{\mathbf{v}}=T_{\text{max}}$. As expected, we do not observe improved performance for $T_{\mathbf{v}}=T_{\text{train}}$ since the transformer itself already performs optimally. The LTV trained with the maximum number of examples is included for comparison purposes, although it does not provide insights into generalizability, as it is exposed to all $T_{\text{max}}$ examples.
Transformer Transformer + LoRA
Transformer + FV (optimized) Transformer + ICV (tuned)
Transformer + LTV
@@ D.1 Complete Set of Evaluations on Regression Tasks
Transformer Transformer + LoRA
Transformer + FV (optimized) Transformer + ICV (tuned)
Transformer + LTV
@@
Transformer Transformer + LoRA
Transformer + FV (optimized) Transformer + ICV (tuned)
Transformer + LTV
@@
D.2 Distributional Shift
We identify two scenarios from [8] where the transformer model’s performance notably degrades during ICL inference: noisy linear regression and skewed covariance matrix.
Noisy linear regression
Noise is added to the output of each example in the form of a standard Gaussian distribution. Specifically, the $i$-th output is defined as $w^{\top}x_{i}+\epsilon_{i}$, where $\epsilon_{i}\sim\mathcal{N}(0,1)$. While the transformer and LTV are trained on standard linear regression, the data during the ICL inference phase is modified by this additive noise. We observe that while the performance of FV considerably degrades, LTV is only slightly perturbed. Specifically, LTV requires training on more examples to maintain the same performance compared to the noise-free setting. For instance, the performance of LTV at $T_{\mathbf{v}}=71$ matches what is achieved in a noise-free environment at around $T_{\mathbf{v}}=56$.
Skewed covariance
The inputs for the prompts are sampled from a zero-mean skewed Gaussian distribution: $x\sim\mathcal{N}(0,\tilde{\Sigma})$, where the eigenbasis of the skewed covariance matrix $\tilde{\Sigma}$ is chosen uniformly at random. Each $i$-th eigenvalue of this matrix is proportional to $1/i^{2}$. The results we obtain align with the expectations based on [8], with the error curves are more unstable and oscillate more than in the isotropic Gaussian case. While LTV also shows some vulnerability to this instability, it still maintains a low mean error and delivers substantial performance enhancement.
@@ D.2 Distributional Shift
D.3 Histograms for Ablation Studies
The histograms effectively illustrate the (unnormalized) probability density functions of the last hidden states. The plots correspond well with the computed KL divergence scores reported in Table 2. As the KL divergence values decrease, the histograms show greater alignment. Specifically, as the training prompt length for the LTV configurations increases, their density profiles become narrower, more closely resembling the shape of the vanilla transformer’s distribution at $T_{\text{train}}$. From a different perspective, this visual alignment supports our hypothesis once again: An optimized LTV with sufficiently long prompts performs near-optimally, as it effectively maintains the last hidden state distribution close to that of the model performing under $T=T_{\text{train}}$.

### v2

@@ C.3.2 Prompting
$\texttt{Q:}\{x_{1}\}\texttt{\textbackslash nA:}\{y_{1}\}\texttt{\textbackslash n}\texttt{\textbackslash n}\ldots\texttt{Q:}\{x_{\text{query}}\}\texttt{\textbackslash nA:}$
@@ C.3.3 Evaluation
We uniformly sample in-context examples (for few-shot) and a query input $N=256$ times, assessing performance as the accuracy if the first token generated by the model matches the first token of the query label (if the label consists of more than one token). To compute the margin of error (denoted after $\pm$ in Tables 1 and 2), we first calculate the standard error of the mean (SEM) as follows:
$\operatorname{SEM}=\frac{\text{standard deviation}}{\sqrt{N}}.$
Next, the t-critical value is determined using the t-distribution with a 95% confidence level and degrees of freedom (N1N-1). The margin of error is then calculated as:
$\text{Margin of error}=t_{\text{critical}}\times\text{SEM}.$
@@ C.4.1 Learning from Regression Data
We recognize that generating distinct functions and prompts at each gradient step could potentially provide an infinite variety of data and functions. This raises concerns about whether LTV overfits to the function classes. Although it was argued that the likelihood of the model encountering a training-similar prompt is extremely low (Garg et al., 2023), we opted for a static dataset approach. Our experiments were conducted using this dataset, with evaluations performed on the prompts constructed from different weight vectors during inference.
@@ C.6 Detailed Ablation Studies
We estimate the probability distributions using Kernel Density Estimation (KDE), employing the standard KDE implementation from scipy (Virtanen et al., 2020) with default settings. However, directly estimating a probability distribution in a high-dimensional space often leads to the curse of dimensionality, where the volume of data required to effectively estimate the distribution grows exponentially with the number of dimensions. A practical solution to this challenge is to employ SVD for dimensionality reduction. This involves decomposing the data matrices $X_{i}\in\mathbb{R}^{M\times d}$ as:
@@ C.7 Computational Resources
Transformer Transformer + LoRA Transformer + FV (optimized)
Transformer + ICV (tuned) Transformer + LTV
@@
Appendix D Complete Set of Results
We present the complete set of results that could not be included in the main body due to page limitation. Specifically, this section includes:
  1. 1.
    Regression Tasks: Evaluations with different LTV training lengths $T_{\mathbf{v}}$
  2. 2.
    Regression Tasks: Evaluations with chain-of-thought prompting based on the findings of Li et al. (2023c)
  3. 3.
    Regression Tasks: Evaluations under distributional shifts
  4. 4.
    NLP Tasks: Accuracy scores (with zero- and few-shot prompting) across 15 NLP benchmarks
  5. 5.
    Ablation Studies: Complete ablation results, including KL divergence scores across different $T_{\mathbf{v}}$ and the corresponding histograms illustrating the distributions of the optimally in-context-learned and LTV-integrated models
D.1 Evaluations on Regression Tasks
The loss curves are shown in Figures 5, 6, and 7. As expected, no performance improvement is observed for $T_{\mathbf{v}}=T_{\text{train}}$, as the transformer already achieves optimal performance. For visual comparison, the LTV trained with the maximum number of examples is included, though it does not provide insights into generalizability, as it is exposed to all $T_{\text{max}}$ examples.
Transformer Transformer + LoRA Transformer + FV (optimized)
Transformer + ICV (tuned) Transformer + LTV
@@ D.1 Evaluations on Regression Tasks
Transformer Transformer + LoRA Transformer + FV (optimized)
Transformer + ICV (tuned) Transformer + LTV
@@
D.2 Chain-of-Thought Prompting
Li et al. (2023c) investigate how CoT enhances ICL for compositional functions, formalizing it as a two-phase process: filtering relevant data and applying ICL on the filtered input. During filtering, the model focuses on the relevant tokens in the prompt based on a given instruction. It then processes the filtered prompt to generate an intermediate output before proceeding to the next step in the sequence.
Multi-layer networks can be viewed as compositional functions since the output depends on the hidden layers, which are functions of the input. Accordingly, Li et al. (2023c) specifically study 2-layer ReLU networks, as in our work. In this context, the relevant tokens at the $k$-th step correspond to the output of the $k$-th hidden layer. For instance, let $z_{k}$ represent the $k$-th intermediate step. During inference, the relevant tokens for $z_{k}$ are $\{z_{k-1}^{i},z_{k}^{i}\}_{i=1}^{T}$, where $T$ is the number of examples in the prompt.
Building on this, we demonstrate that LTV can generalize to capture latent representations at intermediate steps, enabling task fidelity in our length generalization benchmark using Li et al.’s (2023c)’s proposed CoT method. This provides a more flexible and scalable framework for compositional reasoning.
Transformer Transformer + LTV Transformer + CoT-I
Transformer + CoT-I/O Transformer + CoT-I + LTV Transformer + CoT-I/O + LTV
(a) $T_{\mathbf{v}}=126$
(a) $T_{\mathbf{v}}=151$
(a) $T_{\mathbf{v}}=176$
Figure 8: Evaluation on the class of 2-layer ReLU neural networks, with the transformer pre-trained on up to $T_{\text{train}}=101$ chain-of-thought examples per prompt. Results are averaged over a batch of 256 randomly selected tasks. The shaded area represents the 95% confidence interval over the sampled prompts. $T_{\mathbf{v}}$ denotes the prompt length used during LTV training.
D.2.1 Experimental Details
Model Training
As in Appendix C.2.2, we pre-train a custom GPT-2 model (with the same configuration) separately for the two variants of Li et al. (2023c)’s method: CoT-I and CoT-I/O. The task involves 2-layer ReLU network regression, i.e., $f(x)=W_{2}\text{ReLU}(W_{1}x)$. The prompt is structured as $p^{f}=\{x_{1},g(x_{1}),f(x_{1}),\dots,x_{T},g(x_{T}),f(x_{T})\}$, where $g(x)=\text{ReLU}(W_{1}x)$ represents an intermediate computation.
For each $i\leq T$, CoT-I/O performs iterative two-step predictions:
$\hat{s}^{(1)}_{i}=M_{\theta}\big(p_{\mathrm{CoT}}(i)\big),$
$\hat{y}_{i}=\hat{s}^{(2)}_{i}=M_{\theta}\big(\{p_{\mathrm{CoT}}(i),\hat{s}^{(1)}_{i}\}\big)$. (5)
In contrast, CoT-I directly predicts the final output in a single step:
$\hat{y}_{i}=M_{\theta}\big(p_{\mathrm{CoT}}(i)\big).$
The objective functions for pre-training the GPT-2 model are expressed for CoT-I/O and CoT-I respectively as:
[…]
@@ D.2.1 Experimental Details
Optimizing the Learnable Task Vector
In CoT-I/O, we train the LTV for each intermediate step. The objective functions for training LTVs for $g(x)$ and $f(x)$ are, respectively,
[…]
Here, $p_{\text{CoT}}(T_{v}+1)$ consists of $T_{v}$ support examples and a new query:
$\{x_{1},g(x_{1}),f(x_{1}),\dots,x_{T_{v}},g(x_{T_{v}}),f(x_{T_{v}}),x_{\text{query}}\}.$
The LTVs, $v_{\Phi}^{g}=\{v^{g}_{i}\}_{i=1}^{L}$ and $v_{\Phi}^{f}=\{v^{f}_{i}\}_{i=1}^{L}$, are applied at corresponding layers and steps, modifying transformer outputs as $\tilde{M}_{\theta}(\cdot\mid v_{\Phi}^{g})$ and $\tilde{M}_{\theta}(\cdot\mid v_{\Phi}^{f})$.
In CoT-I, a single LTV is trained directly for the final step:
[…]
This formulation allows CoT-I to bypass intermediate step modeling, focusing solely on final output prediction.
We construct the pre-training and LTV datasets following the same setting as in Appendix C.4.1, sampling $N$ i.i.d. tasks from an isotropic Gaussian with $T_{\text{train}}=101$ and $N=100\times 256=25,600$.
D.2.2 Results
Results are provided in Figure 8. For simplicity, we do not include the baselines (i.e., FV and ICV). First, we observe that CoT-I/O (both with and without LTV) performs poorly. This is expected because the final prediction depends on the intermediate step prediction, i.e., $\hat{s}^{(1)}_{i}$ in (5). If the intermediate step introduces noise, it propagates into the final prediction, leading to degraded performance. Consequently, a more direct approach, i.e., CoT-I, achieves better and more effective prediction performance.
The latter case is particularly notable, as the model’s in-context solving abilities improve up to $T=T_{\text{train}}$, showing stable and lower error levels. However, CoT-I still struggles to maintain task fidelity when $T>T_{\text{train}}$ on our benchmark.
For LTV, we observe that whether CoT is used or not, LTV alone can sustain task behavior. This indicates that CoT, when combined with LTV, performs well primarily due to LTV, with minimal contribution to preserving task fidelity. This suggests that our method does not require additional ICL enhancements to maintain task fidelity.
D.3 Distributional Shift
We identify two scenarios from Garg et al. (2023) where the transformer model’s performance notably degrades during ICL inference: noisy linear regression and skewed covariance matrix.
Noisy linear regression
Noise is added to the output of each example in the form of a standard Gaussian distribution. Specifically, the $i$-th output is defined as $w^{\top}x_{i}+\epsilon_{i}$, where $\epsilon_{i}\sim\mathcal{N}(0,1)$. While the transformer and LTV are trained on standard linear regression, the data during the ICL inference phase is modified by this additive noise. In Figure 9, we observe that while the performance of FV degrades significantly, LTV is only slightly affected. Specifically, LTV requires training on more examples to maintain the same performance as in the noise-free setting. For instance, the performance of LTV at $T_{\mathbf{v}}=71$ is comparable to its performance in a noise-free environment at around $T_{\mathbf{v}}=56$.
@@ D.3 Distributional Shift
Dataset Prompt Transformer + FV + ICV + SV + TV + LTV (ours)
Abstractive
AG News (Zhang et al., 2015) Zero-shot 0.0 $\pm$ 0.0 30.1 $\pm$ 5.7 33.6 $\pm$ 5.6 31.7 $\pm$ 5.7 34.5 $\pm$ 5.0 46.9 $\pm$ 6.2
Antonym (Nguyen et al., 2017) Zero-shot 2.3 $\pm$ 1.9 35.2 $\pm$ 5.9 46.4 $\pm$ 5.6 35.2 $\pm$ 6.1 29.6 $\pm$ 6.0 53.5 $\pm$ 6.2
CommonsenseQA (Talmor et al., 2019) Zero-shot 10.2 $\pm$ 3.7 23.0 $\pm$ 5.2 28.9 $\pm$ 5.2 21.7 $\pm$ 1.1 22.0 $\pm$ 4.0 41.0 $\pm$ 6.1
English-French (Lample et al., 2018) Zero-shot 0.0 $\pm$ 0.0 23.4 $\pm$ 5.2 40.2 $\pm$ 5.3 24.4 $\pm$ 2.7 25.4 $\pm$ 5.1 53.5 $\pm$ 6.2
English-Spanish (Lample et al., 2018) Zero-shot 0.0 $\pm$ 0.0 9.0 $\pm$ 3.5 16.5 $\pm$ 4.3 9.7 $\pm$ 3.5 9.4 $\pm$ 2.9 20.7 $\pm$ 5.0
Landmark-Country (Hernandez et al., 2024) Zero-shot 0.0 $\pm$ 0.0 60.9 $\pm$ 6.0 62.7 $\pm$ 5.1 62.9 $\pm$ 5.6 69.9 $\pm$ 6.0 71.1 $\pm$ 5.6
Person-Instrument (Hernandez et al., 2024) Zero-shot 0.0 $\pm$ 0.0 11.3 $\pm$ 3.9 17.1 $\pm$ 4.9 13.1 $\pm$ 2.2 30.1 $\pm$ 3.5 33.2 $\pm$ 5.8
Person-Occupation (Hernandez et al., 2024) Zero-shot 0.0 $\pm$ 0.0 4.3 $\pm$ 2.5 17.2 $\pm$ 3.9 9.0 $\pm$ 2.9 8.1 $\pm$ 5.8 36.3 $\pm$ 5.9
Person-Sport (Hernandez et al., 2024) Zero-shot 0.0 $\pm$ 0.0 2.7 $\pm$ 2.0 12.9 $\pm$ 3.5 12.5 $\pm$ 3.1 12.6 $\pm$ 3.0 16.0 $\pm$ 4.5
Product-Company (Hernandez et al., 2024) Zero-shot 0.0 $\pm$ 0.0 29.3 $\pm$ 5.6 43.0 $\pm$ 5.3 28.2 $\pm$ 4.9 28.7 $\pm$ 5.6 66.4 $\pm$ 5.8
Sentiment Analysis (Socher et al., 2013b) Zero-shot 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 4.8 $\pm$ 1.9 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 10.9 $\pm$ 3.8
Synonym (Nguyen et al., 2017) Zero-shot 1.2 $\pm$ 1.3 2.7 $\pm$ 2.0 14.2 $\pm$ 3.6 2.6 $\pm$ 4.6 2.5 $\pm$ 3.3 18.8 $\pm$ 4.8
Extractive
NER-person (Tjong Kim Sang and De Meulder, 2003) Zero-shot 5.5 $\pm$ 2.8 48.8 $\pm$ 6.2 49.9 $\pm$ 5.8 46.3 $\pm$ 1.3 47.8 $\pm$ 5.3 58.2 $\pm$ 6.1
NER-location (Tjong Kim Sang and De Meulder, 2003) Zero-shot 6.6 $\pm$ 3.1 37.9 $\pm$ 6.0 26.4 $\pm$ 5.1 34.6 $\pm$ 3.9 27.2 $\pm$ 4.7 27.3 $\pm$ 5.5
NER-organization (Tjong Kim Sang and De Meulder, 2003) Zero-shot 21.9 $\pm$ 5.1 48.8 $\pm$ 6.2 52.5 $\pm$ 6.0 46.1 $\pm$ 1.2 46.6 $\pm$ 4.3 64.1 $\pm$ 5.9
Table 2: Complete set of accuracy scores (%) for zero-shot and few-shot (5-shot) predictions, averaged across 256 random seeds. $\pm$ represents the margin of error at a 95% confidence level. The highest accuracy is marked in boldface, and the statistically best-performing method is highlighted.
Transformer 1.098 1.110 0.103
+ Function Vector 0.646 0.572 0.223
+ LTV ($T_{\mathbf{v}}=\{41,41,101\}$) 1.000 1.056 0.095
+ LTV ($T_{\mathbf{v}}=\{42,42,102\}$) 0.952 0.984 0.037
+ LTV ($T_{\mathbf{v}}=\{56,56,126\}$) 0.532 0.260 0.016
+ LTV ($T_{\mathbf{v}}=\{71,71,151\}$) 0.460 0.203 0.017
+ LTV ($T_{\mathbf{v}}=\{86,86,176\}$) 0.344 0.199 0.013
+ LTV ($T_{\mathbf{v}}=\{101,101,201\}$) 0.300 0.096 0.038
D.4 Natural Language Processing
The complete set of accuracy scores on NLP benchmarks is reported in Table 2. Detailed descriptions of these benchmarks can be found in Appendix C.3.1.
D.5 Ablation Studies
The KL divergence scores are reported in Table 3. Additionally, the histograms in Figures 13, 14, and 15 illustrate the (unnormalized) probability density functions of the last hidden states, based on 25,600 collected samples.
The plots correspond well with the computed KL divergence scores. As the KL divergence values decrease, the histograms show greater alignment. Specifically, as the training prompt length for the LTV configurations increases, their density profiles become narrower, more closely resembling the shape of the vanilla transformer’s distribution at $T_{\text{train}}$. From a different perspective, this visual alignment supports our hypothesis once again: An optimized LTV with sufficiently long prompts performs near-optimally, as it effectively maintains the last hidden state distribution close to that of the model performing under $T=T_{\text{train}}$.